OpenStudy (anonymous):

completing the square x^2 - 3x +2 =0

OpenStudy (anonymous):

x^2-2*(3/2)x+(3/2)^2=(3/2)^2-2 (x-3/2)^2=9/4-2 (X-3/2)^2=1/4 x-3/2=+-1/2 x=2,1

OpenStudy (lalaly):

\[\large{(x-\frac{3}{2})^2-(\frac{3}{2})^2+2=0}\]\[\large{(x-\frac{3}{2})^2-\frac{9}{4}+\frac{8}{4}}=0\]\[\large{(x-\frac{3}{2})^2=\frac{1}{4}}\]\[\large{x-\frac{3}{2}= \pm \frac{1}{2}}\]\[x=\frac{1}{2}+\frac{3}{2}=2\]\[x=\frac{-1}{2}+\frac{3}{2} = 1\]